OpenStudy (anonymous):

prove that 3:8 is equivalent to 12:32. a.use diagrams to support your answer. b.use the description of equivalent ratios to support your answer.

OpenStudy (anonymous):

they are all factors of four

OpenStudy (mathmale):

prove that 3:8 is equivalent to 12:32: I'd suggest that you write the ratio 12:32 as\[\frac{ 12 }{ 32}\] and then factor both the 12 and the 32. Write out the factors:\[\frac{ 12 }{ 32 }=\frac{ ( ? ) ( ? ) }{ ( ? ) ( ? ) }\] and then reduce this last result by cancelling out the common factor.

OpenStudy (anonymous):

thanks!